# Patherea: Cell Detection and Classification for the 2020s

### v1

@@ Abstract
This paper presents a Patherea, a framework for point-based cell detection and classification that provides a complete solution for developing and evaluating state-of-the-art approaches. We introduce a large-scale dataset collected to directly replicate a clinical workflow for Ki-67 proliferation index estimation and use it to develop an efficient point-based approach that directly predicts point-based predictions, without the need for intermediate representations. The proposed approach effectively utilizes point proposal candidates with the hybrid Hungarian matching strategy and a flexible architecture that enables the usage of various backbones and (pre)training strategies. We report state-of-the-art results on existing public datasets - Lizard, BRCA-M2C, BCData, and the newly proposed Patherea dataset. We show that the performance on existing public datasets is saturated and that the newly proposed Patherea dataset represents a significantly harder challenge for the recently proposed approaches. We also demonstrate the effectiveness of recently proposed pathology foundational models that our proposed approach can natively utilize and benefit from. We also revisit the evaluation protocol that is used in the broader field of cell detection and classification and identify the erroneous calculation of performance metrics. Patherea provides a benchmarking utility that addresses the identified issues and enables a fair comparison of different approaches. The dataset and the code will be publicly released upon acceptance.
@@
Automating cell detection and classification represents a key step for performing diagnostics in a more efficient and reproducible manner and has been one of the first topics to be addressed in histopathology image analysis Xing and Yang (2016). Supervised-learning-based approaches require expertly labeled datasets, which are costly to obtain in sufficient quantities in comparison with natural images in the broader computer vision domain. Most approaches utilize point-based annotations, which are easier to obtain in comparison with bounding boxes or segmentation masks and are most often not needed to support the end application in terms of diagnostics. Various point-based deep-learning-based approaches have been proposed in recent years (Xie et al., 2015, 2018; Lee et al., 2021; Abousamra et al., 2021; Huang et al., 2023b, a; Pina et al., 2024) that require the use of intermediate representations to regress and classify cell centers, or the use of existing complex DETR-based (Carion et al., 2020) object-detection architectures, specialized to support point annotations. In contrast, we provide a direct point-to-point approach, that utilizes general computer vision backbones (Dosovitskiy et al., 2021; Liu et al., 2022), with simple architecture, fast convergence, and support for the use of existing large-scale pathology pretrained foundational models.
Generating labeled data for different possible diseases is not only prohibitively expensive and time-consuming, but in the medical domain, often impossible due to low occurrence and natural and technical variability present in the samples. On the other hand, diagnostics is performed on a daily basis around the world, with workflows being digitized and unlabeled samples being stored in large quantities. Similar to general language and vision domains (Radford et al., 2019, 2021; He et al., 2022; Oquab et al., 2024), pathology foundational models have been recently presented (Filiot et al., 2023; Chen et al., 2024; Vorontsov et al., 2023; Dippel et al., 2024; Xu et al., 2024; Nechaev et al., 2024) that utilize vast amounts of unlabeled data from up to millions of samples from various diseases, stainings and instruments.
Limited labeled training data is needed to train or fine-tune the models on a specific downstream task of interest, as well as to benchmark different approaches. Publicly available cell detection and classification datasets in pathology are rare and mostly consists of the standard hematoxylin and eosin (H&E) staining and with at most couple of tens of thousands annotated cell centers, often annotated with an semi-automatic approach and in a patch-based manner, which does not reflect the actual pathologists workflow (Sirinukunwattana et al., 2016; Huang et al., 2020; Graham et al., 2021; Abousamra et al., 2021; Ryu et al., 2023). Different metrics and their implementations - some of them flawed, are used, making it difficult to compare different proposed approaches.
@@
We propose a novel end-to-end point-based cell detection and classification architecture based on Vision Transformers that completely removes the need for pre-processing and post-processing, enabling fast convergence and utilization of domain-specific foundational models.
@@
To the best of our knowledge, we present111TBD the largest manually labeled point-based Ki-67 dataset that was collected by directly mimicking the clinical practice, utilizing the full-resolution WSIs. We demonstrate that the performance on existing public datasets is saturated and that the newly proposed dataset represents a significantly harder challenge for the recently proposed approaches.
@@
We evaluate the proposed approaches on existing public and newly proposed cell detection and classification datasets and improve upon recently proposed state-of-the-art approaches - in some cases by a significant margin.
@@
We make all of our Patherea code (Patherea-P2P, Patherea-FCRN, benchmarking code) publicly available for research use222TBD.
@@ 2.1 Point-based Cell Detection
Our proposed Patherea-P2P approach (Section 3) builds upon P2PNet (Song et al., 2021) and introduces a Hybrid Hungarian Matching to increase the amount of positive supervision in comparison with the vanilla one-to-one matching utilized in P2PNet, as well as DETR-based approaches. Implicitly, this also incorporates contextual information. We also design our architecture around a standard visual backbone with a lightweight head, in comparison with object detection specific DETR-based architectures, which limit the use of large pre-trained foundational models.
@@ 2.2 Datasets
Cell (nuclei) detection and classification requires much more fine-grained labels. A selection of the most prominent datasets is presented in Table 1.
@@
OCELOT (Ryu et al., 2023) are more recent datasets that were specifically designed for the task of point-based cell detection and classification tasks. BCData (Huang et al., 2020) represents a similar task to our proposed Patherea-Breast (Section 4.2) dataset, and the only dataset that uses stains beyond the standard H&E, but was collected using a standard microscope, with an attached camera. OCELOT (Ryu et al., 2023) additionally provides a larger field-of-view pair of cell and tissue annotations.
@@
2.3 Pathology Foundational Models
@@
Similar to the general computer vision domain, pathology foundational models have been recently proposed (Filiot et al., 2023; Chen et al., 2024; Vorontsov et al., 2023; Dippel et al., 2024; Xu et al., 2024; Nechaev et al., 2024) that utilize vast amounts of digitized pathology data from various organs, stains and instruments. Most of the models utilize the standard H&E data, which is available in abundance. Recent foundational models (Dippel et al., 2024; Nechaev et al., 2024) have also incorporated different stains. Most of the models are trained on proprietary training data and are not publicly available. They are predominately evaluated on various downstream clinical tasks, where a significant boost is observed in comparison with ImageNet pre-trained models (Campanella et al., 2024). We utilize a recently provided open-source Hibou pathology foundational model (Nechaev et al., 2024), based on the ViT-B backbone and apply it to the task of point-based cell detection and classification. This presents a novel application of an existing foundational model to the specific downstream task of point-based cell detection and classification.
@@ 3 Methods
The proposed Patherea-P2P architecture is presented in Figure 1. The overall architecture was inspired by the crowd counting and localization Point-to-Point Network (P2PNet) (Song et al., 2021), which was further extended with a novel Hybrid Hungarian-based regression loss, classification capability and modern building blocks that enable the use of self-supervised approaches to train or use existing foundational models tailored for digital pathology. The used notation mostly follows (Song et al., 2021).
@@ 3.1 P2P Network
Patherea-P2P method training input represents a (histology tissue) patch with a set of point-based annotations , which represents a cell center location and a cell type for a particular cell - among labeled cells and cell types in a given patch (Figure 1, left, green dots). The trained model similarly predicts a set of points in a given patch , based on the proposal candidates (Figure 1, top-right, red dots), with an additional confidence score for each of the possible cell types, which we threshold to produce a final set of predictions (Figure 1, bottom-right, red dots). Predicted points should be as close as possible to ground-truth cell locations and of the same type , while maximizing a confidence score , for the task of cell localization. Similarly, the overall number of predicted number of cells () should be as close as possible to the actual number of cells in a given patch, for the task of cell counting. Patherea-P2P combines and optimizes both tasks at the same time, while taking into account individual cell types.
@@
Let denote the optimal permutation , such that represents the matched proposal candidate for ground truth , while represents unmatched proposal candidates that are treated as background. The classification loss can thus be derived as following (2):
@@ 3.2 Hybrid Hungarian Matching
3.3 Foundational Backbone
The proposed Patherea-P2P method removes the need for any intermediate representations on the input side, as well as any post-processing, thus making it fully end-to-end. Additionally, we design the architecture in a simplified manner, reducing the need for complex task-specific building blocks, beyond the light-weight heads. Most of the existing work (Huang et al., 2023b, a; Pina et al., 2024) base their architecture on DETR-based frameworks. Vision, or even tasks-specific (e.g., object detection) approaches are usually more complex, but have usually achieved better performance due to the inherent inductive bias present in the architecture. We design the architecture around general-purpose architectures like ConvNext (Liu et al., 2022) and ViT (Dosovitskiy et al., 2021), which directly support (multi-modal) pre-training in a self-supervised manner. We utilized ConvNext and ViT backbones, together with Feature Pyramid Networks (FPN) (Lin et al., 2017), which enables to use of higher-resolution features, beneficial for dense prediction tasks. For ViT, we additionally utilized ViT-Adapter (Chen et al., 2023), which allows plain ViT to achieve comparable performance to vision-specific transformers.
Masked image modelling can be used to train foundational models from scratch using MAE (He et al., 2022) for ViT or SparK (Tian et al., 2023) for ConvNext. Alternatively, existing, open-source pathology foundational models (Nechaev et al., 2024) can be used, which are most often only available for ViTs.
@@ 4.1 Background
Ki-67 proliferation index is calculated as a proportion of positive tumor cells against the negative tumor cells. For example, International Ki-67 in Breast Cancer Working Group (Dowsett et al., 2011) recommends counting at least 1000 tumor cells in at least three high-power (x40 objective) fields. In practice, "eye-balling" is often used (Reid et al., 2015), which basically means estimating the percentage of Ki-67 positive tumor cells by "scanning" the entire slide at intermediate power (x10 objective), without actually counting the individual cells. This results in a low inter-observer agreement (Reid et al., 2015) in comparison with manual counting. The development of AI-based approaches holds the potential to enable scoring Ki-67 at a fraction of the time needed, even in comparison with eye-balling, while reaching the inter-observer agreement of the manual counting. The focus of this work is to develop an efficient approach for cell detection and classification and compare its performance against the ground truth data from pathologists. Clinical relevance of an AI-assisted Ki-67 proliferation scoring and comparison against the pathologists will be the scope of the subsequent publications.
4.2 Dataset
@@
In comparison with the related work (Sirinukunwattana et al., 2016; Huang et al., 2020; Graham et al., 2021; Abousamra et al., 2021; Ryu et al., 2023), we directly used WSIs for labeling and not the selected extracted patches, such that we directly mimicked the actual clinical workflow, if it were to be digitized. We deployed the Digital Slide Archive platform (Gutman et al., 2017) in a cloud environment for the management of WSIs, as well as for the labeling. The selected slides were labeled by 4 pathologists and each of the pathologists had access to the platform and labeled the slides through a web browser. The pathologists were free to select the regions to be labeled, which were marked with a polygon, with all of the cells and their types labeled - as depicted in Figure 2. The pathologists had access to the WSI and were able to freely navigate across the slide and use different magnifications to label the cells. Five different classes of cells were labeled: positive and negative tumor cells, cells that show proliferation, but are not tumor cells - others positive, others negative, as well as normal healthy cells. The selected samples represent biopsies of the cancer tissue, resulting in a low occurrence of healthy cells. As described in Section 4.1, the ratio between positive and negative tumor cells is important to derive the Ki-67 proliferation index.
We divided the samples across three different sets - LNET, GNET and Breast. LNET part of the dataset represents the NET samples from larynx, pharynx and lungs. GNET part of the dataset represents the NET samples from the gastrointestinal tract. The Breast part of the dataset represents breast cancer samples. Each sample was labeled by one senior pathologist. Publicly available datasets are usually labeled by junior pathologists or students and later verified by a senior pathologist. In comparison, all of our samples were labeled by senior pathologists with an average of around 25 years of working experience in clinical practice. Some large-scale datasets (Graham et al., 2021) are labeled in a semi-automatic fashion, where cell centers are first proposed by an automatic (e.g. AI) approach, which are then verified by a pathologist. In comparison, we labeled all of our samples completely manually, thus producing high-quality annotations.
@@
| LNET | 20 | 25,118 | 35,288 | 660 | 10,334 | 80 | 71,480 |
| GNET | 22 | 6,458 | 43,606 | 761 | 10,762 | - | 61,587 |
| Breast-P1 | 29 | 9,665 | 17,187 | 908 | 8,913 | - | 36,673 |
| Breast-P2 | 29 | 8,813 | 18,583 | 274 | 5,477 | - | 33,147 |
| 100 | 50,054 | 114,664 | 2,603 | 35,486 | 80 | 202,887 |
Nevertheless, there is some inherent bias with assigning cell types, including with senior pathologists. This is especially evident with border cases, such as other positive and negative cells. We have thus labeled the Breast part of the Patherea dataset by two pathologists. The same regions that were selected by the first pathologist were given to the second pathologist, which labeled all of the cells in the regions that were selected by the first pathologist. This is depicted in Table 2 by Breast-P1 and Breast-P2.
Annotated WSIs were then patched into patches of and split into 3 folds, that were used to train and evaluate different models in Section 6. The dataset, including the raw WSIs and the patched folds, will be made publicly available.
@@ 5 Evaluation Protocol
The Hungarian algorithm is used by most of the established public datasets and benchmarks (Xie et al., 2018; Graham et al., 2019; Huang et al., 2023b, a) to perform the one-to-one matching between the detections and labeled ground truth data. Some of the established benchmarks (Graham et al., 2019) have performed the Hungarian matching in a flawed manner333https://github.com/vqdang/hover_net/blob/master/metrics/stats_utils.py#L393, which does not report the correct F1 scoring. Such flawed scoring is then used by the following work (Huang et al., 2023b, a), which prevents a fair comparison against different approaches and datasets, if not the same benchmarking code is used - which would still result in an underestimated F1 score. The problem arises due to the global cost optimization nature of the Hungarian algorithm, if directly a distance matrix is used to do the assignments. This is depicted in Figure 3 - right, that depicts the flawed nature of assignments.
@@
Due to the global objective optimization, a distant detection, correctly detected for one of the ground truth cells is assigned to another cell, that has no close detection (e.g., Figure 3, left arrow). A detection, close-by, but outside of the ground truth circular region is assigned to the first ground truth location (e.g., Figure 3, right arrow). After the Hungarian matching, filtering based on the ground truth radius is used. Due to the wrongly predicted point being matched to the first cell, an additional false-negative is produced, despite the first cell being correctly detected in the first place. This is the result of the global nature of the Hungarian optimization objective.
The solution, correctly implemented in the broader domain of crowd counting and localization (Wang et al., 2020) is to perform the Hungarian algorithm on a Boolean matrix, obtained by thresholding the distance matrix based on the ground truth radius . This ensures that the matching is correct, with the only limitation being that there is no guarantee that if multiple detections are present in a circular region, the closest one to the center will be selected. This does not influence the correctness of the reported F1 score.
@@
The alternative greedy approach of iterating across all the predicted points and ground truth data is used by some of the related work (Abousamra et al., 2021), which is also often flawed444https://github.com/TopoXLab/MCSpatNet/blob/main/03_eval_localization_fscore.py#L17, by allowing multiple predictions within the ground truth radius count as true-positives (one-to-many matching), which overestimates the F1 score.
The benchmarking code will be released alongside the dataset presented in Section 4 that addresses the above-mentioned problems to enable a reproducible and robust comparison across different datasets and approaches.

### v2

@@ Abstract
We present Patherea, a unified framework for point-based cell detection and classification that enables the development and fair evaluation of state-of-the-art methods. To support this, we introduce a large-scale dataset that replicates the clinical workflow for Ki-67 proliferation index estimation. Our method directly predicts cell locations and classes without relying on intermediate representations. It incorporates a hybrid Hungarian matching strategy for accurate point assignment and supports flexible backbones and training regimes, including recent pathology foundation models. Patherea achieves state-of-the-art performance on public datasets—Lizard, BRCA-M2C, and BCData—while highlighting performance saturation on these benchmarks. In contrast, our newly proposed Patherea dataset presents a significantly more challenging benchmark. Additionally, we identify and correct common errors in current evaluation protocols and provide an updated benchmarking utility for standardized assessment. The Patherea dataset and code are publicly available to facilitate further research and fair comparisons.
@@
Automated cell detection and classification are essential for efficient and reproducible diagnostics, and they form a fundamental topic in histopathology image analysis Xing and Yang (2016). Supervised learning approaches depend on expertly labeled datasets, which are expensive and time-consuming to obtain at scale compared to natural image datasets in computer vision. Most methods rely on point-based annotations, which are more accessible than bounding boxes or segmentation masks and typically sufficient for diagnostic applications. Recent deep learning approaches (Xie et al., 2015, 2018; Lee et al., 2021; Abousamra et al., 2021; Huang et al., 2023b, a; Pina et al., 2024) employ intermediate representations to regress and classify cell centers or adapt complex DETR-based (Carion et al., 2020) object detection architectures for point annotations. In contrast, our approach provides direct point-to-point mapping using standard computer vision backbones (Dosovitskiy et al., 2021; Liu et al., 2022), offering a simpler architecture with faster convergence and compatibility with existing large-scale pathology foundation models.
Generating labeled data for different possible diseases is not only prohibitively expensive and time-consuming, but in the medical domain, often impossible due to low occurrence and natural and technical variability present in the samples. On the other hand, diagnostics is performed daily around the world, with workflows being digitized and unlabeled samples being stored in large quantities. Similar to general language and vision domains (Radford et al., 2019, 2021; He et al., 2022; Oquab et al., 2024), pathology foundation models have been recently presented (Filiot et al., 2023; Chen et al., 2024; Vorontsov et al., 2023; Dippel et al., 2024; Xu et al., 2024; Nechaev et al., 2024) that utilize vast amounts of unlabeled data from up to millions of samples from various diseases, stainings and instruments.
Limited labeled training data is essential both for fine-tuning models on specific downstream tasks and for benchmarking different approaches. However, publicly available datasets for cell detection and classification in pathology are scarce. Most existing datasets are based on standard hematoxylin and eosin (H&E) staining and contain, at most, a few tens of thousands of annotated cell centers. These annotations are often produced semi-automatically and limited to small image patches, which does not reflect the real-world workflow of pathologists (Sirinukunwattana et al., 2016; Huang et al., 2020; Graham et al., 2021; Abousamra et al., 2021; Ryu et al., 2023). Furthermore, the use of inconsistent evaluation metrics and implementation details across studies makes it challenging to compare the performance of different methods in a meaningful way.
@@
We propose a novel end-to-end point-based cell detection and classification architecture based on Vision Transformers that completely removes the need for pre-processing and post-processing, enabling fast convergence and utilization of domain-specific foundation models.
@@
We present the largest manually labeled point-based Ki-67 dataset, annotated by four senior pathologists on full-resolution WSIs following clinical protocols. Our dataset reveals performance saturation on existing public datasets and poses significantly greater challenges for current state-of-the-art approaches.
@@
We demonstrate significant improvements over state-of-the-art methods on both existing and newly proposed cell detection and classification datasets, achieving superior F1 scores while maintaining faster training and inference speeds.
@@
We make our Patherea framework (dataset, code) publicly available for research use111https://github.com/patherea/patherea-p2p.
@@ 2.1 Point-based Cell Detection
Point-based cell detection can be viewed as a simplified alternative to instance segmentation methods, which have become prevalent in the pathology domain (Graham et al., 2019, 2023; Hörst et al., 2024, 2025). While instance segmentation provides detailed cellular boundaries, it relies on labor-intensive pixel-level annotations that are often impractical to obtain at scale. In contrast, point annotations are significantly cheaper and faster to acquire. Moreover, for many downstream pathology tasks–such as cell counting, phenotyping, or spatial analysis—the exact cell shape offers limited additional utility beyond the cell’s location. As such, point-based methods offer a more annotation-efficient solution while still meeting the requirements of many pathology applications.
Our proposed Patherea-P2P approach (Section 3) builds upon P2PNet (Song et al., 2021) and introduces a Hybrid Hungarian Matching to increase the amount of positive supervision compared to the vanilla one-to-one matching utilized in P2PNet and DETR-based approaches. Implicitly, this also incorporates contextual information. We also design our architecture around a standard visual backbone with a lightweight head, in comparison with object detection specific DETR-based architectures, which limit the use of large pre-trained foundation models.
@@ 2.2 Datasets
Cell (nuclei) detection and classification requires much finer-grained labels. A selection of the most prominent datasets is presented in Table 1.
@@
OCELOT (Ryu et al., 2023) are more recent datasets that were specifically designed for the task of point-based cell detection and classification tasks. BCData (Huang et al., 2020) represents a similar task to our proposed Patherea-Breast dataset (Section 4), and the only dataset that uses stains beyond the standard H&E, but was collected using a standard microscope, with an attached camera. OCELOT additionally provides a larger field-of-view pair of cell and tissue annotations (Ryu et al., 2023).
@@
2.3 Pathology Foundation Models
@@
Similar to the general computer vision domain, pathology foundation models have been recently proposed (Filiot et al., 2023; Chen et al., 2024; Vorontsov et al., 2023; Dippel et al., 2024; Xu et al., 2024; Nechaev et al., 2024) that utilize vast amounts of digitized pathology data from various organs, stains and instruments. Most of the models utilize the standard H&E data, which is available in abundance. Recent foundation models (Dippel et al., 2024; Nechaev et al., 2024) have also incorporated different stains. Most of the models are trained on proprietary training data and are not publicly available. They are predominantly evaluated on various downstream clinical tasks, where a significant boost is observed in comparison with ImageNet pre-trained models (Campanella et al., 2024). Task-specific approaches that utilize foundation models have also started to appear—such as recently presented CellViT++ (Hörst et al., 2025) instance segmentation-based approach that reduces the need for segmentation masks in training.
We utilize a recently provided open-source Hibou pathology foundation model (Nechaev et al., 2024), based on the ViT-B backbone and apply it to the task of point-based cell detection and classification. This presents a novel application of an existing foundation model to the specific downstream task of point-based cell detection and classification.
@@ 3 Methods
The proposed Patherea-P2P architecture is presented in Figure 1. The overall architecture was inspired by the crowd counting and localization Point-to-Point Network (P2PNet) (Song et al., 2021), which was further extended with a novel Hybrid Hungarian-based regression loss, classification capability and modern building blocks that enable the use of self-supervised approaches to train or use existing foundation models tailored for digital pathology. The used notation mostly follows P2PNet (Song et al., 2021).
@@ 3.1 P2P Network
Patherea-P2P method training input represents a (histology tissue) patch with a set of point-based annotations , which represents a cell center location and a cell type for a particular cell - among labeled cells and cell types in a given patch (Figure 1, left, green dots for one particular cell class ). The trained model similarly predicts a set of points in a given patch , based on the proposal candidates (Figure 1, top-right, red dots), with an additional confidence score for each of the possible cell types, which we threshold to produce a final set of predictions (Figure 1, bottom-right, red dots). Predicted points should be as close as possible to ground-truth cell locations and of the same type , while maximizing a confidence score , for the task of cell localization. Similarly, the overall number of predicted number of cells () should be as close as possible to the actual number of cells in a given patch, for the task of cell counting. Patherea-P2P combines and optimizes both tasks at the same time, while taking into account individual cell types.
@@
Let denote the optimal permutation , such that represents the matched proposal candidate for ground truth , while represents unmatched proposal candidates that are treated as background. The classification loss can thus be derived as follows (2):
@@ 3.2 Hybrid Hungarian Matching
3.3 Foundation Model Support
The proposed Patherea-P2P method removes the need for any intermediate representations on the input side, as well as any post-processing, thus making it fully end-to-end. Additionally, we design the architecture in a simplified manner, reducing the need for complex task-specific building blocks, beyond the light-weight heads. Most of the existing work (Huang et al., 2023b, a; Pina et al., 2024) base their architecture on DETR-based frameworks. Vision, or even task-specific (e.g., object detection) approaches are usually more complex, but have usually achieved better performance due to the inherent inductive bias present in the architecture. We design the architecture around general-purpose architectures like ConvNext (Liu et al., 2022) and ViT (Dosovitskiy et al., 2021), which directly support (multi-modal) pre-training in a self-supervised manner. We utilized ConvNext and ViT backbones, together with Feature Pyramid Networks (FPN) (Lin et al., 2017), which enables to use of higher-resolution features, beneficial for dense prediction tasks. For ViT, we additionally utilized ViT-Adapter (Chen et al., 2023), which allows plain ViT to achieve comparable performance to vision-specific transformers.
Masked image modelling can be used to train foundation models from scratch using MAE (He et al., 2022) for ViT or SparK (Tian et al., 2023) for ConvNext. Alternatively, existing, open-source pathology foundation models (Nechaev et al., 2024; Chen et al., 2024) can be used, which are most often only available for ViTs.
@@ 4 Patherea Dataset
To ensure clinical realism, we utilized whole slide images (WSIs) for annotation rather than pre-selected image patches, closely mimicking a digitized clinical workflow. The Digital Slide Archive (DSA) platform (Gutman et al., 2017) was deployed in a cloud environment for managing and labeling WSIs, accessible to pathologists via a standard web browser.
Four pathologists participated in the annotation process. Each had unrestricted access to the entire WSI, allowing them to freely navigate, zoom across magnifications, and select regions of interest for annotation. This flexibility enabled the selection of more representative and diverse tissue areas, reducing sampling bias compared to fixed patch-based approaches. Annotated regions were outlined with polygons and all visible cells within these regions were exhaustively labeled with their respective cell types (Figure 4).
Four cell categories were defined: (i) positive tumor cells, (ii) negative tumor cells, (iii) positive non-tumor cells, and (iv) negative non-tumor cells. This labeling protocol reflects both diagnostic relevance and the clinical decision-making process (A).
We organized samples into three anatomically distinct groups: LNET (larynx, pharynx, and lung NET samples), GNET (gastrointestinal tract NET samples), and Breast (breast cancer samples). All annotations were performed by senior pathologists with an average of 25 years of clinical experience. In contrast to publicly available datasets that typically rely on junior pathologists or students with senior verification, our labeling approach ensures consistent annotation quality. Unlike large-scale datasets that employ semi-automatic labeling approaches (Graham et al., 2021), where AI-proposed labels are subsequently verified, our dataset features entirely manual annotations, providing high-quality ground truth labels.
@@
| LNET | 20 | 25,118 | 35,288 | 660 | 10,414 | 71,400 |
| GNET | 22 | 6,458 | 43,606 | 761 | 10,762 | 61,587 |
| Breast-P1 | 29 | 9,665 | 17,187 | 908 | 8,913 | 36,673 |
| Breast-P2 | 29 | 8,813 | 18,583 | 274 | 5,477 | 33,147 |
| 100 | 50,054 | 114,664 | 2,603 | 35,486 | 202,807 |
Nevertheless, there is some inherent bias with assigning cell types, including with senior pathologists. This is especially evident with borderline cases, such as other positive and negative cells. We have thus labeled the Breast part of the Patherea dataset by two pathologists. The same regions that were selected by the first pathologist were given to the second pathologist, which labeled all of the cells in the regions that were selected by the first pathologist. This is depicted in Table 2 by Breast-P1 and Breast-P2.
Annotated WSIs were then divided into patches and split into 3 folds for model training and evaluation in Section 6. The dataset and fold splits are made publicly available222https://huggingface.co/datasets/ds2268/patherea.
@@ 5 Evaluation Protocol
The Hungarian algorithm is used by most of the established public datasets and benchmarks (Xie et al., 2018; Graham et al., 2019; Huang et al., 2023b, a) to perform the one-to-one matching between the detections and labeled ground truth data. Some of the established benchmarks (Graham et al., 2019) have performed the Hungarian matching incorrectly333https://github.com/vqdang/hover_net/blob/master/metrics/stats_utils.py#L393, which does not report the correct F1 scoring. Such incorrect scoring is then used by the following work (Huang et al., 2023b, a), which prevents a fair comparison against different approaches and datasets, if not the same benchmarking code is used - which would still result in an underestimated F1 score. The problem arises due to the global cost optimization nature of the Hungarian algorithm, if a distance matrix is used directly to do the assignments. This is depicted in Figure 2 - right, which depicts the flawed nature of assignments.
@@
Due to the global objective optimization, a distant detection, correctly detected for one of the ground truth cells (Figure 2, C) is assigned to another cell that has no close detection (Figure 2, A). A detection, close-by, but outside of the ground truth circular region is assigned to the first ground truth location (Figure 2, B). After the Hungarian matching, filtering based on the ground truth radius is used. Due to the wrongly predicted point being matched to the first cell (Figure 2, C), an additional false-negative is produced, despite the first cell being correctly detected in the first place.
This is the result of the global nature of the Hungarian optimization objective, which minimizes the total cost across all assignments (Equation 7):
where is the weight (distance) between predicted point and ground truth point , and is a binary variable indicating assignment. A detection is correct if it falls within a specified radius of its corresponding ground truth. However, the Hungarian algorithm’s global optimization can assign a detection to a distant ground truth.
The solution, correctly implemented in the broader domain of crowd counting and localization (Wang et al., 2020), involves applying the Hungarian algorithm to a Boolean matrix obtained by thresholding a distance matrix based on the ground truth radius . Given two point sets of predicted points and ground truth points , we first construct a weighted bipartite graph , where edge weights represent the Euclidean distance between each predicted point and each ground truth point . We then threshold this graph at radius , retaining only edges with weights to produce a Boolean match matrix. The Hungarian algorithm is applied to this thresholded graph to obtain the maximum bipartite matching, from which we compute true-positives (TP), false-positives (FP), and false-negatives (FN). We experimentally confirm the difference between the two approaches in B.4.
@@
The alternative greedy approach of iterating across all the predicted points and ground truth data is used by some of the related work (Abousamra et al., 2021), which is also incorrect444https://github.com/TopoXLab/MCSpatNet/blob/main/03_eval_localization_fscore.py#L17, by allowing multiple predictions within the ground truth radius count as true-positives (one-to-many matching), which overestimates the F1 score.